OpenStudy (melodious):

which is smaller 5 root 10 or 4 root 9?????????

OpenStudy (anonymous):

You mean \(5\sqrt{10}\) or \(4\sqrt{9}\)? Observe that \(5>4\) and \(\sqrt{10}>\sqrt{9}\). Which do you think is larger?

OpenStudy (melodious):

4 root 9 is larger right?

OpenStudy (anonymous):

No. Since 4<5 and sqrt9 < sqrt10, one would expect that 4sqrt9<5sqrt10, otherwise one would violate some variant of Archimedes' principle.

OpenStudy (melodious):

can u simplify the roots and help me out @nettle404

OpenStudy (anonymous):

I think \(\sqrt{10}\) is about as simple as it can get. On the other hand, what squared gives you \(9\)?

OpenStudy (melodious):

3

OpenStudy (anonymous):

Exactly, so \(\sqrt9=3\). But \(\sqrt{10}\) has no simple expression.

OpenStudy (melodious):

but the question is 4 root 9 so it means 4 multiplied by root 9

OpenStudy (melodious):

which is 3x4=12

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

You got it.

OpenStudy (melodious):

what about 5 root 10?

OpenStudy (anonymous):

Not easily simplified. Can you think of any way to square a number to obtain 10?

zepdrix (zepdrix):

You can introduce another number like this if it helps you understand what is going on. \[\large\rm 4\sqrt9\lt5\sqrt9\]But also\[\large\rm 5\sqrt9\lt5\sqrt{10}\]So we see that\[\large\rm 4\sqrt{9}\lt5\sqrt9\lt5\sqrt{10}\]Therefore\[\large\rm 4\sqrt9\lt5\sqrt{10}\]

zepdrix (zepdrix):

Maybe that's more complicated though :3 lol

OpenStudy (anonymous):

How about proof by contradiction? Suppose that \(5\sqrt{10}\leq4\sqrt{9}\), then \(5\sqrt{10}\leq12\) and \(\sqrt{10}\leq12/5\). Since \(\sqrt{10}>3\) but \(12/5<3\), we have \(3<3\), which is false.